Hi
the ratio of adams age to bobs age is 7:3.adam is 84 years old.in how many years time will their total
age be 126.
my answer is 6 textbook says 3
thank you

Answer by rothauserc(4718)

let x be Bob's age
:
7/3 = 84/x
:
cross multiply the fractions
:
7x = 252
:
x = 36
:
Bob is 36 years old
:
Their current total is 120, let y be "in how many years"
:
(84+y) + (36+y) = 126
:
2y + 120 = 126
:
2y = 6
:
y = 3
:
***********************************************
In 3 years the total of their ages will be 126
***********************************************
: